Fill in the data table below for a projectile that is fired off of a building 57° BELOW the horizontal with a speed of 46 m/sHow do I get the acceleration? also what is the formula i need to use to find the other velocity values?
Answer #1

Initial velocity of projectile u=46\,m/s

As the projectile is fired below horizontal with angle \theta=57\degree

Initial horizontal velocity is U u cos 0 = 46 cos(-57°) 46 cos(-57°) = 25.053 m

Initial vertical velocity is u_y=u\sin\theta=46\sin(-57\degree)=-38.579\,m/s

As the projectile moves in the air, the horizontal component of velocity does not change with time v_x=u_x=25.053\,m/s as there is no acceleration is the horizontal direction a_x=0\,m/s^2 ( at all times) and vertical component of velocity changes with time v_y=u_y-gt as the acceleration in the vertical direction is a_y=-g=-9.81\,m/s^2 (at all times)

At time t=0\,s

v_y=-38.579-9.81*0=-38.579\,m/s and a_y=-9.81\,m/s^2

At time t=1\,s

v_y=-38.579-9.81*1=-48.389\,m/s and a_y=-9.81\,m/s^2

At time t=2\,s

v_y=-38.579-9.81*2=-58.199\,m/s  and a_y=-9.81\,m/s^2 ​​​​​​​

At time t=3\,s

v_y=-38.579-9.81*3=-68.009\,m/s  and a_y=-9.81\,m/s^2 ​​​​​​​

At time t=4\,s

v_y=-38.579-9.81*4=-77.819\,m/s and a_y=-9.81\,m/s^2 ​​​​​​​

t(s) 0 1 2 3 4
vx (m/s) 25.053 25.053 25.053 25.053 25.053
ax(m/s2 ) 0 0 0 0 0
vy(m/s) -38.579 -48.389 -58.199 -68.009 -77.819
ay(m/s2) -9.81 -9.81 -9.81 -9.81 -9.81